OpenStudy (anonymous):

A simple pendulum has a period of 15 seconds. What is the length of the pendulum in meters? I got 55.9, I am not sure if this is correct.

OpenStudy (anonymous):

T=2pi sqrt (l/g) T^2 = 4*pi^2 (l/g) l=(g*T^2)/4pi^2 i am getting 57.04 did you take g=9.8m/s^2